# Bistability and chaos in the discrete two-gene Andrecut-Kauffman model

### v1

@@ 8. A search for positive Lyapunov exponents in both systems
Figure 7. Pairs of parameters $\beta$ and $\varepsilon$ for the models (1) and (2) for which there exists a parameter $\alpha\in[0,100]$, or a pair of parameters $(\alpha_{1},\alpha_{2})\in[0,100]\times[0,100]$, respectively, with a positive maximum Lapunov exponent. Pairs with a positive maximum Lyapunov exponent observed in both models are marked in blue, while those with this observation made only for model 2 are marked in orange.
@@ 9. The maximum Lyapunov exponent for various parameters $\alpha_{1}$ and $\alpha_{2}$
Plots of Lyapunov exponents as a function of the parameters of a system convey detailed information on the changes in the behavior of the system when the parameters vary. In Figure 8, we show the maximum Lyapunov exponent computed for the model (2) when $\alpha_{1}$ and $\alpha_{2}$ vary within the interval $(0,250)$ while $\beta=0.2$, $\varepsilon=0.6$, and $n=3$ are fixed.
Due to the color scale chosen in Figure 8, regions of stability are shown in black (low values of the maximum Lyapunov exponent), whereas chaotic dynamics is indicated by lighter shades of gray, with brightness corresponding to the magnitude of chaos, measured by the value of the maximum Lyapunov exponent. As one can see in the figure, moving across the parameter space along a straight line might involve multiple transitions between chaos and stability. Indeed, the bifurcation diagram in Figure 9 was computed for the parameters along the line segment plotted in red in Figure 8. It shows how chaos arises in the system from a series of period doubling bifurcations at small values of $\alpha_{1}$ and $\alpha_{2}$. One can also see the next major stable region that appears when the chaotic attractor collapses into a period-five orbit.
Figure 8. The maximum Lyapunov exponent computed for the model (2) for $\alpha_{1},\alpha_{2}\in(0,250)$ with $\beta=0.2$, $\varepsilon=0.6$, and $n=3$.
Figure 9. Bifurcation diagram of the $x$ variable computed along the segment in the parameter space plotted in red in Figure 8.
@@ 10. Asymmetry in the computation of the maximum Lyapunov exponent
The symmetry in the equations defining this dynamical system implies that, for a global attractor, Lyapunov exponent values on the $(\alpha_{1},\alpha_{2})$ parameter plane should exhibit symmetry around the $\alpha_{1}=\alpha_{2}$ line, as demonstrated in Figure 8. However, while scanning the parameter space using a fixed set of $16$ initial conditions, we identified regions where that symmetry was unexpectedly broken. This phenomenon can be explained by the presence of multiple attractors in the phase space.
In order to illustrate this, let us consider a fixed initial condition $x_{0}$, and let $A_{1}$ and $A_{2}$ be two attractors of the system that are present for certain values of $\alpha_{1}$ and $\alpha_{2}$. Assume $x_{0}$ lies within the basin of attraction $B_{1}$ associated with $A_{1}$. Now consider the system with the parameters $\alpha_{1}$ and $\alpha_{2}$ swapped, that is, $\alpha_{1}^{\prime}=\alpha_{2}$ and $\alpha_{2}^{\prime}=\alpha_{1}$. The new system will behave identically to the original one, except the phase space will be reflected by the $y=x$ line. Consequently, attractors $A_{1}^{\prime}=T(A_{1})$ and $A_{2}^{\prime}=T(A_{2})$ will now exist, where $T\colon(x,y)\mapsto(y,x)$. Their basins of attraction will similarly be reflections $B_{1}^{\prime}=T(B_{1})$, $B_{2}^{\prime}=T(B_{2})$ of the original ones, and thus it is now possible for $x_{0}$ to lie within $B_{2}^{\prime}$. While the attractors’ positions have changed, their underlying properties, including their Lyapunov spectra, remain the same.
Therefore, if the maximal Lyapunov exponent was $\lambda_{1}$ in $A_{1}$ and $\lambda_{2}$ in $A_{2}$, with $\lambda_{1}\neq\lambda_{2}$ the maximal Lyapunov exponent calculated at the point $x_{0}$ may differ when the parameters $\alpha_{1}$ and $\alpha_{2}$ are swapped, despite the intrinsic symmetry of the equations that define the model.
This behavior can be utilized for detecting ranges of parameters where forms of bistability occur. In the following section we focus on the effects of changing $\varepsilon$. We sampled $1024$ uniformly spaced values of $\varepsilon$ from $[0,1]$. For each sample, the maximum Lyapunov exponent was calculated for $(\alpha_{1},\alpha_{2})$ in a $512\times 512$ grid over the range of $(0,250)$, resulting in a square array of exponents denoted $L$. The amount of asymmetry $S^{\prime}$ was then calculated as
@@
Figure 10. Asymmetry of the system assessed for $\alpha_{1},\alpha_{2}\in(0,250)$, $\beta=0.2$ and $n=3$ measured by the two quantities introduced in the text: $S^{\prime}$ shown as the black line and $p^{\pm}$ indicated by the shaded area.
Two intervals with high asymmetry values can be noticed. While the results for $\varepsilon\in(0,0.16)$ appear to be caused by slow convergence of the system with low but nonzero coupling between variables, making them inconclusive, the method converges relatively quickly for $\varepsilon\in(0.7,1)$. Indeed, closer analysis of parameters from the latter subset reveal the presence of bistability that is discussed in the next section.
11. Three kinds of bistability in the system
For wide ranges of parameters, both models (1) and (2) exhibit the phenomenon of bistability. It is defined as the existence of two coexisting attractors, sometimes with different characteristics, such as a periodic orbit or a chaotic attractor. Any trajectory in the system then converges to one attractor or the other one, which is reflected in the system experiencing a different type of long-term behavior, depending on the initial conditions.
@@ 11. Three kinds of bistability in the system
Table 1. Examples of parameters resulting in different modes of bistability of the model (2).
@@
Figure 11. An attracting cycle (blue) and a chaotic attractor (orange) observed in the system (2) with the parameters $\alpha_{1}=43.6$, $\alpha_{2}=75.7$, $\varepsilon=0.9$, $\beta=0.2$, and $n=3$ mentioned in Table 1. Approximations of attraction basins of the attractors are indicated in the corresponding bright colors.
To the best of our knowledge, this phenomenon was not previously observed for such wide ranges of parameters in the Andrecut-Kauffman model. Up to now, the only kind of bistability known for this system was the two stable fixed points present for $\varepsilon=1$ (see [AK06]).
A possible reason for the fact that the bistability of this kind was not discovered in this system by previous studies is that while multiple attractors exist for $\alpha_{1}=\alpha_{2}$, their maximum Lyapunov exponents are identical and therefore cannot be detected by this feature alone. Therefore, admitting different rates $\alpha_{1}\neq\alpha_{2}$ in the model allows better understanding of possible dynamics of gene expression system.
Finally, it turns out that for $\varepsilon=0$, the existence of more than two attractors is possible. The variables $x$ and $y$ are independent in that case, and each of them might have its own attractor. In particular, at some values of the parameters, both may have a stable period-4 cycle; let us denote these cycles by $(x_{1},x_{2},x_{3},x_{4})$ and $(y_{1},y_{2},y_{3},y_{4})$, respectively. Then the cartesian product of the two 4-element sets is an attractor for the two-dimensional system. It splits into four disjoint attracting cycles:
@@ 12. Conclusions and future research directions
Gene expression is a fundamental biological process that influences nearly every aspect of the functioning of cells, including cell differentiation, responses to stimuli, and cell cycle regulation. Research on the dynamics of gene expression provides insight into mechanisms that lead to stable and chaotic patterns of activity, which is crucial in both systems biology and synthetic biology, where the goal is to design controlled genetic systems with specific properties.
@@
Our results illustrate different types of dynamical behavior encountered across different parameter regimes. These results shed light on the influence of parameters on system’s stability and its potential for exhibiting chaotic behavior. In particular, the existence of system configurations with the chaos-order type of bistability shown in Section 11 has profound implications for the real gene expression system. A sudden external impulse (eg. an injection containing one of the transcription factors) or continuous addition of one protein into the system can permanently destabilize the system by pushing it across the boundary between basins of attraction towards the chaotic region. In the same way, ordered state can be reestablished by applying some force towards the stable attractor until the system switches its state. Similar switching between stable states could also occur naturally during processes analogous to cell specialization or cells working in different modes.
Further directions of research on the model (2) might include the analysis of the dynamics for $n>4$.
One might also be interested in finding unstable fixed points or unstable periodic orbits, especially in the case of bistability.
Thorough analysis of the actual bifurcations that take place in the system when the parameters change might shed light on the actual role of the various parameters on the dynamics exhibited by the system and on the stability of solutions.
@@
Finally, comprehensive analysis of global dynamics, taking into account all the possible values of parameters in the ranges indicated in Section 3, focused on detecting interesting phenomena like bistability or chaotic dynamics, might provide information on how often these phenomena occur and how likely they are to be encountered in a real system. For example, one could try applying the approach introduced in [AKK+09] and further developed in [PSRG23] for this purpose, combined with additional analyses, such as computation of the maximum Lyapunov exponent discussed in Section 7.
@@ Acknowledgments
The authors M. Rosman and A. Bartłomiejczyk acknowledge support obtained from Gdańsk University of Technology in the framework of the DEC-27/1/2022/IDUB/III.4c/Tc grant under the Technetium Talent Management Grants – ‘Excellence Initiative – Research University’ program. The authors M. Palczewski and P. Pilarczyk acknowledge support received from the National Science Centre, Poland, within the grant OPUS no. 2021/41/B/ST1/00405.

### v2

@@ 8. A search for positive Lyapunov exponents in both models
Figure 7. Pairs of parameters $\beta$ and $\varepsilon$ for models (1) and (2) for which there exists a parameter $\alpha\in[0,100]$, or a pair of parameters $(\alpha_{1},\alpha_{2})\in[0,100]\times[0,100]$, respectively, with the positive maximum Lapunov exponent. Pairs with the positive maximum Lyapunov exponent observed in both models are marked in blue, while those with this observation made only for model (2) are marked in orange.
@@ 9. The maximum Lyapunov exponent for various parameters $\alpha_{1}$ and $\alpha_{2}$
Plots of Lyapunov exponents as a function of parameters of a model convey detailed information on the changes in the behavior of the model when the parameters vary. In Figure 8, we show the maximum Lyapunov exponent computed for model (2) when $\alpha_{1}$ and $\alpha_{2}$ vary within the interval $(0,250)$ while $\beta=0.2$, $\varepsilon=0.6$, and $n=3$ are fixed.
Due to the color scale chosen in Figure 8, stability regions are shown in black (low values of the maximum Lyapunov exponent), and chaotic dynamics is indicated by lighter shades of gray, with brightness corresponding to the magnitude of chaos, measured by the value of the maximum Lyapunov exponent. As one can see in the figure, moving across the parameter space along a straight line might involve multiple transitions between chaos and stability. Indeed, the bifurcation diagram in Figure 9 was computed for the parameters taken along the line segment plotted in red in Figure 8. It shows how chaos arises in the model from a series of period-doubling bifurcations at small values of $\alpha_{1}$ and $\alpha_{2}$. One can also see the next major stable region that appears when the chaotic attractor collapses into a period-five orbit.
Figure 8. The maximum Lyapunov exponent computed for model (2) for $\alpha_{1},\alpha_{2}\in(0,250)$ with $\beta=0.2$, $\varepsilon=0.6$, and $n=3$.
Figure 9. Bifurcation diagram of the $x$ variable computed along the segment in the parameter space plotted in red in the previous figure (Figure 8), parametrized as $(\alpha_{1},\alpha_{2})=(20+130t,10+215t)$ for $t\in[0,1]$.
@@ 10. Asymmetry in the computation of the maximum Lyapunov exponent
The symmetry in the equations that define this dynamical system implies that, for a global attractor, Lyapunov exponent values in the $(\alpha_{1},\alpha_{2})$ parameter plane should exhibit symmetry around the $\alpha_{1}=\alpha_{2}$ line, as demonstrated in Figure 8. However, while scanning the parameter space using a fixed set of $16$ initial conditions, we identified regions where that symmetry was unexpectedly broken. This phenomenon can be explained by the presence of multiple attractors in the phase space.
In order to illustrate this phenomenon, we consider a fixed initial condition $x_{0}$, and let $A_{1}$ and $A_{2}$ be two attractors of the model that are present for certain values of $\alpha_{1}$ and $\alpha_{2}$. Assume $x_{0}$ is in the basin of attraction $B_{1}$ associated with $A_{1}$. Now consider the model with the parameters $\alpha_{1}$ and $\alpha_{2}$ swapped, that is, $\alpha_{1}^{\prime}=\alpha_{2}$ and $\alpha_{2}^{\prime}=\alpha_{1}$. The new model will behave identically to the original one, except that the phase space will be reflected by the $y=x$ line. Consequently, attractors $A_{1}^{\prime}=T(A_{1})$ and $A_{2}^{\prime}=T(A_{2})$ will now exist, where $T\colon(x,y)\mapsto(y,x)$. Their basins of attraction will similarly be reflections $B_{1}^{\prime}=T(B_{1})$, $B_{2}^{\prime}=T(B_{2})$ of the original ones, and thus it is now possible for $x_{0}$ to lie within $B_{2}^{\prime}$. Although the attractors’ locations have changed, their underlying properties, including their Lyapunov spectra, remain the same.
Therefore, if the maximum Lyapunov exponent was $\lambda_{1}$ in $A_{1}$ and $\lambda_{2}$ in $A_{2}$, with $\lambda_{1}\neq\lambda_{2}$ the maximum Lyapunov exponent calculated at the point $x_{0}$ may differ when the parameters $\alpha_{1}$ and $\alpha_{2}$ are swapped, despite the intrinsic symmetry of the equations that define the model.
This behavior can be utilized to detect ranges of parameters where certain forms of bistability occur. The following computations focused on the effects of choosing $\varepsilon$. We sampled $1024$ uniformly spaced values of $\varepsilon$ from $[0,1]$. For each sample, the maximum Lyapunov exponent was calculated for $(\alpha_{1},\alpha_{2})$ in a $512\times 512$ grid over the range of $(0,250)$, resulting in a square array of exponents denoted by $L$. The amount of asymmetry $S^{\prime}$ was then calculated as
@@
Figure 10. Asymmetry of the model assessed for $\alpha_{1},\alpha_{2}\in(0,250)$, $\beta=0.2$ and $n=3$ measured by the two quantities introduced in the text: $S^{\prime}$ shown as the black line and $p^{\pm}$ indicated by the shaded area.
Two intervals with high asymmetry values can be noticed. The results for $\varepsilon\in(0,0.16)$ appear to reflect slow convergence of the model with a low but nonzero coupling between the variables, and are thus inconclusive. However, the method converges relatively quickly for $\varepsilon\in(0.7,1)$. Indeed, closer analysis of parameters from the latter subset reveals the presence of bistability which we discuss in the next section.
11. Three kinds of bistability in the model
For wide ranges of parameters, both models (1) and (2) exhibit the phenomenon of bistability. It is defined as the existence of two coexisting attractors, sometimes with different characteristics, such as a periodic orbit or a chaotic attractor. Any trajectory in the model then converges to one attractor or to the other one, which is reflected in the model experiencing a different type of long-term behavior, depending on the initial conditions.
@@ 11. Three kinds of bistability in the model
Table 1. Examples of parameters resulting in different modes of bistability of model (2).
@@
Figure 11. An attracting cycle (blue) and a chaotic attractor (orange) observed in model (2) with the parameters $\alpha_{1}=43.6$, $\alpha_{2}=75.7$, $\varepsilon=0.9$, $\beta=0.2$, and $n=3$ mentioned in Table 1. Approximations of attraction basins of the attractors are indicated by the corresponding bright colors.
To the best of our knowledge, this phenomenon was not previously observed for such wide ranges of parameters in the Andrecut-Kauffman model. Up to now, the only kind of bistability known for this model was the two stable fixed points present for $\varepsilon=1$ (see [4]).
A possible reason for the fact that the bistability of this kind was not discovered in this model by previous studies is that while multiple attractors exist for $\alpha_{1}=\alpha_{2}$, their maximum Lyapunov exponents are identical and therefore cannot be detected by this feature alone. Therefore, admitting different rates $\alpha_{1}\neq\alpha_{2}$ in the model allows for better understanding of possible dynamics of the actual gene expression system.
Finally, it turns out that for $\varepsilon=0$, the existence of more than two attractors is possible. The variables $x$ and $y$ are independent in that case, and each of them might have its own attractor. In particular, at some values of the parameters, both may have a stable period-4 cycle; let us denote these cycles by $(x_{1},x_{2},x_{3},x_{4})$ and $(y_{1},y_{2},y_{3},y_{4})$, respectively. Then the Cartesian product of the two 4-element sets is an attractor for the two-dimensional model. It splits into four disjoint attracting cycles:
@@ 12. Conclusions and future research directions
Gene expression is a fundamental biological process that influences nearly every aspect of cell function, including cell differentiation, responses to stimuli, and cell cycle regulation. Research on the dynamics of gene expression provides insight into mechanisms that lead to stable and chaotic patterns of activity, which is crucial in both systems biology and synthetic biology, where the goal is to design controlled genetic systems with specific properties.
@@
Our results illustrate different types of dynamical behavior encountered across different parameter regimes. These results shed light on the influence of parameters on stability of the model, and thus on the modeled system, and its potential for exhibiting chaotic behavior as well. In particular, the existence of configurations with the chaos-order type of bistability shown in Section 11 has profound implications for the real gene expression system. A sudden external impulse (e.g., an injection containing one of the transcription factors) or continuous addition of one protein into the system can permanently destabilize the system by pushing it across the boundary between basins of attraction towards the chaotic region. In the same way, ordered state can be reestablished by applying some force towards the stable attractor until the system switches its state. Similar switching between stable states could also occur naturally during processes analogous to cell specialization or cells working in different modes.
The process of multimerization is essential for the functionality of many proteins, as they achieve full functionality only in their multimeric forms, such as dimers ($n=2$), trimers ($n=3$) or higher-order multimers, depending on cellular conditions. Therefore, there is a natural demand for further research on model (2) to include the analysis of the dynamics also for $n>4$. As one can see in Figure 7, there are more parameters that exhibit chaotic dynamics for $n=4$ than for $n=3$. Therefore, one can expect that the analysis of dynamics becomes more demanding with the increase in $n$.
Another promising direction of further research might be to search for unstable fixed points and unstable periodic orbits, especially in the case of bistability. Although biological systems typically settle down in stable states, one can apply mild control (e.g., in terms of some external force or an agent) in order to keep the system in an unstable state at minimal cost. Therefore, determining unstable fixed points and unstable periodic solutions is also important for a better understanding of such systems and handling them.
In general, the analysis of bifurcations and unstable fixed points in these types of models is of particular importance, as it allows the identification of conditions that lead to a qualitative change in the behavior of the system, such as transitions between activated and silenced states of genes. Thorough analysis of the actual bifurcations that take place in the system when the parameters change might shed light on the actual role of the various parameters for the dynamics exhibited by the system and on the stability of solutions. As a result of a bifurcation, some stable states might lose their stability. This means that for the gene expression process, a gene that was previously active can become silenced, potentially leading to changes in cellular behavior or function. On the other hand, a new stable solution might emerge. This could correspond to the appearance of a new biological state or pathway that was not previously observed, such as the activation of alternative gene expression mechanisms or regulatory networks. It might also be the case that chaotic solutions emerge, which means that due to the change of some conditions, such as the increase in temperature that affects the intensity of some processes, the system suddenly switches its state from a well-understood stable equilibrium to hard-to-predict chaotic oscillations. From the perspective of gene expression, such chaotic behavior could result in irregular production of proteins, leading to unpredictable cellular responses and potentially impacting the organism’s overall functionality. Both silencing and overexpression as defects in gene regulation can cause profound and unfavorable change in cellular functioning; see, e.g., [13, 14]. As a result, many kinds of bifurcations are important from the biological point of view and should not be neglected.
@@
Finally, comprehensive analysis of global dynamics that takes into account all the possible values of parameters in the ranges indicated in Section 3, focused on detecting interesting phenomena like bistability or chaotic dynamics, might provide information on how often these phenomena occur and how likely they are to be encountered in the model. For example, one could try applying the approach introduced in [6] and further developed in [18] for this purpose, combined with additional analyses, such as the calculation of the maximum Lyapunov exponent discussed in Section 7. Although this approach is limited to the analysis of dynamics at finite resolution, it can be argued that phenomena occurring below certain scale in a mathematical model of a biological system are not biologically relevant; see, e.g., the discussion in [15].
@@ Acknowledgments
The authors M. Rosman and A. Bartłomiejczyk acknowledge support obtained from Gdańsk University of Technology in the framework of the DEC-27/1/2022/ IDUB/III.4c/Tc grant under the Technetium Talent Management Grants – ‘Excellence Initiative – Research University’ program. The authors M. Palczewski and P. Pilarczyk acknowledge support received from the National Science Centre, Poland, within the grant OPUS 2021/41/B/ST1/00405.